# Impartial Games: a Challenge for Reinforcement Learning

### v5

@@ Abstract
AlphaZero-style reinforcement learning (RL) algorithms have achieved superhuman performance in many complex board games such as Chess, Shogi, and Go. However, we showcase that these algorithms encounter significant and fundamental challenges when applied to impartial games, a class where players share game pieces and optimal strategy often relies on abstract mathematical principles. Specifically, we utilize the game of Nim as a concrete and illustrative case study to reveal critical limitations of AlphaZero-style and similar self-play RL algorithms. We introduce a novel conceptual framework distinguishing between champion and expert mastery to evaluate RL agent performance. Our findings reveal that while AlphaZero-style agents can achieve champion-level play on very small Nim boards, their learning progression severely degrades as the board size increases. This difficulty stems not merely from complex data distributions or noisy labels, but from a deeper representational bottleneck: the inherent struggle of generic neural networks to implicitly learn abstract, non-associative functions like parity, which are crucial for optimal play in impartial games. This limitation causes a critical breakdown in the positive feedback loop essential for self-play RL, preventing effective learning beyond rote memorization of frequently observed states. These results align with broader concerns regarding AlphaZero-style algorithms’ vulnerability to adversarial attacks, highlighting their inability to truly master all legal game states. Our work underscores that simple hyperparameter adjustments are insufficient to overcome these challenges, establishing a crucial foundation for the development of fundamentally novel algorithmic approaches, potentially involving neuro-symbolic or meta-learning paradigms, to bridge the gap towards true expert-level AI in combinatorial games.
keywords:
reinforcement learning, impartial games, learning parties, neural network
@@ 1 Introduction
The advent of AlphaZero and its variations has marked a significant success in machine learning, demonstrating superhuman capabilities across complex strategy games such as chess, Go, and shogi [35, 36, 30, 28]. These self-play reinforcement learning (RL) algorithms, leveraging powerful computational resources, can achieve superhuman performance within hours. However, despite these remarkable successes, AlphaZero-style algorithms exhibit inherent limitations, particularly concerning their robustness and generalization to all possible game states. For instance, RL agents like KataGo have been shown to be vulnerable to adversarial attacks [41] and perturbations [18], and may even overlook optimal moves in specific game states, indicating an incomplete mastery of the game’s entire legal state space [42]. The identified vulnerabilities and struggles with specific game structures highlight critical gaps in the generalization capabilities and theoretical understanding of these advanced RL systems. Addressing these limitations is paramount for advancing robust and truly general artificial intelligence, especially as complex decision-making systems become more pervasive.
@@
This paper investigates a distinct class of games called impartial games, where AlphaZero-style algorithms encounter fundamental challenges. Unlike partisan games (e.g., chess, Go), impartial games involve players sharing game pieces, often requiring the implicit learning of a parity function to determine optimal moves. The Nim game, a canonical example of an impartial game, prominently features such a requirement. According to the Sprague–Grundy theorem, every impartial game can be reduced to an equivalent Nim-heap [4], establishing Nim’s central role in impartial game theory. This makes impartial games an ideal testbed for probing the generalization boundaries of current RL paradigms.
@@
Learning parity functions, especially from a uniform data distribution, is a known challenge for neural networks [1, 31, 13, 12, 26, 33, 17]. While self-play RL agents generate non-uniform game positions, potentially creating a curriculum learning environment, empirical evidence suggests that the inherent ’noisy label’ problem in self-play RL significantly impedes neural networks’ ability to model parity functions [43]. Specifically, it has been shown that a modest rate of incorrect labels can degrade parity function learning to random guessing for bitstrings of moderate length.
@@
Building on these theoretical considerations, particularly the statistical neutrality of parity functions [37], we hypothesize that AlphaZero-style algorithms, without specialized modifications, will struggle to master Nim and other impartial games on sufficiently large boards. This challenge primarily stems from the difficulty neural networks face in modeling parity functions, rather than solely from algorithmic limitations. Overcoming this requires either explicit incorporation of combinatorial game theory principles or the development of more robust neural network architectures capable of discerning such complex parity-based game properties.
@@
To explore this hypothesis and gain deeper insights into the challenges faced by RL agents in mastering Nim-like games, we developed and extensively analyzed an AlphaZero-style algorithm tailored for Nim. Our investigation also involved revisiting AlphaZero’s performance in chess, focusing on positions where parity-related problems are crucial for identifying optimal moves. This comprehensive exploration culminated in a detailed theoretical and practical analysis of issues related to impartial games and parity problems. We introduce two novel metrics for evaluating self-play RL agents’ performance—’champion’ and ’expert’ perspectives—which offer a nuanced assessment of mastery beyond conventional winning rates.
@@
Our findings reveal that AlphaZero-style algorithms exhibit significant difficulties in effectively learning to play Nim, especially as the board size increases. Despite Nim’s low inherent complexity and simple optimal strategy from both human and computational perspectives, the algorithms struggle. While the algorithm can achieve proficiency on smaller boards or towards endgames through exhaustive search, larger, non-trivial board positions frequently resulted in the policy network failing to provide valuable guidance to the Monte Carlo Tree Search (MCTS), and the value network performing no better than random guessing. These scenarios represent ’blind spots’ where optimal actions are entirely overlooked [18, 42].
@@
In games like chess, AlphaZero’s Policy-Value Neural Networks (PV-NNs) are imperfect, but this is often compensated by the MCTS. For some Nim boards, this imperfection might not be critical, as the policy network can guide the winning player towards positions that do not demand precise evaluations. However, we contend that this advantage diminishes on sufficiently large Nim boards due to the PV-NNs’ inherent inability to model parity functions, leading to two critical issues:
@@
The remainder of this paper is structured as follows: Section 2 introduces the Nim game and its relationship to other impartial games via the Sprague-Grundy theorem. In Section 3, we revisit AlphaZero and Leela Chess Zero (LCZero) to analyze their strengths and subtle weaknesses, particularly regarding parity-related issues impacting their PV-NN. Section 4 defines our proposed ’champion’ and ’expert’ mastery levels with illustrative examples. Section 5 presents an overview of the AlphaZero-style algorithms and our re-implementation  [^1] The code for the experiments in this paper is publicly available at: https://github.com/sagebei/Impartial-games-RL. We analyse the performance of the trained algorithm from the champion and expert perspective respectively, and then examine the moves made on some Nim board positions, demonstrating the challenges the AlphaZero-style algorithm faces in becoming an expert agent on large boards. We conclude the paper with general remarks, conjectures and directions for further research in Section 6.
2 Impartial Games and Nim
@@
The class of impartial games constitutes a significant subclass of combinatorial games [4, 5, 6, 7]. This category encompasses a diverse range of games, including classic examples such as Take-and-break, Subtraction, Heap, and Poset games. It also includes mathematically intriguing games like Nim, Sprouts, Treblecross, Cutcake, Guiles, Wyt Queens, Kayles, Grundy’s Game, Quarto, Cram, Chomp, Subtract a Square, and Notakto [4, 5]. Many of these impartial games feature multiple variants, adding to their complexity and research interest.
@@
An impartial game is a two-player game characterized by identical sets of available moves from any given position, irrespective of whose turn it is. Players take alternating turns. A player loses if they cannot make a legal move on their turn, meaning the player who makes the last legal move wins. Critically, any position within an impartial game can be definitively classified as either a losing (P-position, previous player winning) or winning (N-position, next player winning) position, indicating whether the current player has no winning move or at least one winning move, respectively.
@@
Nim serves as an important example of an impartial game and is often studied within mathematical game theory due to its elegant and analytically derived solution. Nim is played by two players who take turns removing counters from distinct heaps (or piles) arranged in a row [8, 24]. On each turn, a player must remove at least one counter and may remove any number of counters, provided they are all drawn from a single heap. The objective is to be the player who removes the last counter, thereby leaving an empty board for the opponent. An initial Nim board can be represented as an array of heap sizes: $[n_{1},n_{2},\ldots,n_{k}]$, where each $n_{k}\in\{1,3,\dots,2k-1\}$ typically denotes the initial size of the $k$-th heap. The maximum number of legal moves from an initial position $[n_{1},n_{2},\ldots,n_{k}]$ is $\sum_{j=1}^{k}n_{j}$, as each move requires removing at least one counter. A game position can be represented as an array of numbers $[v_{1},v_{2},\ldots,v_{k}]$, where $v_{j}\leq n_{j}$. Although Nim is often studied in abstract terms with arbitrary configurations, RL algorithms require a fixed board size, and each self-play game begins from a predefined initial state. Figure 1 visually illustrates an initial Nim board, an intermediate game state, and the terminal winning condition.
@@
Figure 1: An illustration of Nim game states. The left panel shows an initial board configuration of five heaps: $[n_{1},n_{2},n_{3},n_{4},n_{5}]=[1,3,5,7,9]$. The middle panel depicts an intermediate board state during gameplay: $[v_{1},v_{2},v_{3},v_{4},v_{5}]=[1,2,4,4,3]$, resulting from players removing counters. The right panel represents the terminal state where all counters are cleared, signifying a win for the player who made the last move.
@@
A property of Nim is that for any given position, it is mathematically straightforward to determine the winning player and identify all available winning moves. This determination is achieved by calculating the binary digital sum of the number of counters in each heap, a process known as the nim-sum [8]. This operation involves summing the heap sizes in binary without carrying over bits. The computational complexity of calculating the nim-sum has been shown to be linear in the number of heaps and logarithmic in their sizes [15], requiring only logarithmic space memory [9]. Consequently, determining whether a given Nim position is a winning or losing state is computationally efficient and deterministic. This analytical solvability stands in stark contrast to the challenges faced by reinforcement learning algorithms in games lacking such perfect information or straightforward winning criteria.
@@
where $N(s)$ represents the set of all states $s^{\prime}$ reachable from $s$ in one legal move, and the mex (minimum excluded value) function returns the smallest non-negative integer not present in a given set [3]. A position with Sprague-Grundy value $G(s)$ behaves identically to a Nim heap of size $G(s)$. Crucially, a position is a losing position if and only if its Sprague-Grundy value is 0. This theorem implies that the insights gained from studying Nim’s learnability extend naturally to a wide array of other impartial games.
The theoretical underpinnings of impartial games are closely connected to the nim-sum, which in turn relies on the concept of the parity function. Thus, the parity function plays an implicit or explicit central role in the theory of impartial games, as many such games can be shown to mimic Nim or parts of Nim. To illustrate this fundamental connection, consider the impartial game Sprouts [4, 16], a relatively simple game invented by John Conway and Michael Paterson. Positions in Sprouts typically yield nimber values of 0, 1, 2, or 3 [6]. As depicted in Figure 2(a), a specific Sprouts position can have a nimber value of 3. Furthermore, a modified Sprouts position (Figure 2(b)), designed as an "isolated land" or "gadget" that cannot interact with external elements, also possesses a nimber value of 3. A Sprouts starting position comprising $n$ copies of the gadget shown in Figure 2(b) effectively mimics any Nim position reachable from a starting position with $n$ heaps, each containing 3 counters.
@@
(b) A Sprouts position designed as an isolated gadget, also having a nimber value of 3. This gadget can be used to construct larger impartial games that behave like Nim (adapted from p. 599 in [6]).
@@
Figure 2: Mimicking Nim with Sprouts gadgets. A game of Nim played on a board [3,3,,3][3,3$\ldots$,3][ 3 , 3 , … , 3 ] with $n$ heaps is equivalent to a Sprouts game constructed with $n$ copies of the isolated gadget shown in Figure 2(b).
@@
Despite the analytical simplicity of Nim, some impartial games present significant computational challenges, as their complexity can rise to PSPACE-complete. From an algorithmic perspective, while many impartial games share the linear-time solvability of Nim, certain others, such as Node Kayles and Geography, are demonstrably harder, being PSPACE-complete [29]. Moreover, even some partisan games, where move legality depends on the player, can exhibit properties akin to impartial games, such as Dawson’s chess, which is effectively an impartial game in disguise [4].
@@
Historically, algorithms for impartial games have predominantly relied on handcrafted programs incorporating principles similar to alpha-beta search, but specifically tailored for the mex operation and Sprague-Grundy values [40]. For instance, [3] introduced a novel method for pruning search trees based on node values computed via the mex function, particularly effective for relatively "short" impartial games like Nim, Chomp, and Cram. However, this approach typically does not scale efficiently to games with large board sizes or extensive state spaces. These conventional, handcrafted programs, which leverage the explicit mathematical structure of impartial games, serve as crucial benchmarks against which the performance and generalization capabilities of emerging reinforcement learning-based algorithms can be rigorously evaluated. The objective is to assess whether contemporary RL methods can not only match but potentially surpass the limitations of these specialized traditional approaches, particularly for larger or more complex instances of impartial games where explicit solution might be intractable.

### v6

@@ Abstract
AlphaZero-style reinforcement learning (RL) algorithms have achieved superhuman performance in many complex board games such as Chess, Shogi, and Go. However, we showcase that these algorithms encounter significant and fundamental challenges when applied to impartial games, a class where players share game pieces and optimal strategy often relies on abstract mathematical principles. Specifically, we utilise the game of Nim as a concrete and illustrative case study to reveal critical limitations of AlphaZero-style and similar self-play RL algorithms. We introduce a novel conceptual framework distinguishing between champion and expert mastery to evaluate RL agent performance. Our findings reveal that while AlphaZero-style agents can achieve champion-level play on very small Nim boards, their learning progression severely degrades as the board size increases. This difficulty stems not merely from complex data distributions or noisy labels, but from a deeper representational bottleneck: the inherent struggle of generic neural networks to implicitly learn abstract, non-associative functions like parity, which are crucial for optimal play in impartial games. This limitation causes a critical breakdown in the positive feedback loop essential for self-play RL, preventing effective learning beyond rote memorisation of frequently observed states. These results align with broader concerns regarding AlphaZero-style algorithms’ vulnerability to adversarial attacks, highlighting their inability to truly master all legal game states. Our work underscores that simple hyperparameter adjustments are insufficient to overcome these challenges, establishing a crucial foundation for the development of fundamentally novel algorithmic approaches, potentially involving neuro-symbolic or meta-learning paradigms, to bridge the gap towards true expert-level AI in combinatorial games.
@@ 1 Introduction
The advent of AlphaZero and its variations has marked a significant success in machine learning, demonstrating superhuman capabilities across complex strategy games such as chess, Go, and shogi (Silver et al., 2017, 2018; Schrittwieser et al., 2020; Sadler and Regan, 2019). These self-play reinforcement learning (RL) algorithms, leveraging powerful computational resources, can achieve superhuman performance within hours. However, despite these remarkable successes, AlphaZero-style algorithms exhibit inherent limitations, particularly concerning their robustness and generalisation to all possible game states. For instance, RL agents like KataGo have been shown to be vulnerable to adversarial attacks (Wang et al., 2023) and perturbations (Lan et al., 2022), and may even overlook optimal moves in specific game states, indicating an incomplete mastery of the game’s entire legal state space (Wu, 2019). The identified vulnerabilities and struggles with specific game structures highlight critical gaps in the generalisation capabilities and theoretical understanding of these advanced RL systems. Addressing these limitations is paramount for advancing robust and truly general artificial intelligence, especially as complex decision-making systems become more pervasive.
@@
This paper investigates a distinct class of games called impartial games, where AlphaZero-style algorithms encounter fundamental challenges. Unlike partisan games (e.g., chess, Go), impartial games involve players sharing game pieces, often requiring the implicit learning of a parity function to determine optimal moves. The Nim game, a canonical example of an impartial game, prominently features such a requirement. According to the Sprague–Grundy theorem, every impartial game can be reduced to an equivalent Nim-heap (Berlekamp et al., 2001), establishing Nim’s central role in impartial game theory. This makes impartial games an ideal testbed for probing the generalisation boundaries of current RL paradigms.
@@
Learning parity functions, especially from a uniform data distribution, is a known challenge for neural networks (Abbe et al., 2023; Shalev-Shwartz et al., 2017; Daniely and Malach, 2020; Cornacchia and Mossel, 2023; Raz, 2018; Shoshani and Shamir, 2025; Han and Ghoshdastidar, 2025). While self-play RL agents generate non-uniform game positions, potentially creating a curriculum learning environment, empirical evidence suggests that the inherent ’noisy label’ problem in self-play RL significantly impedes neural networks’ ability to model parity functions (Zhou and Riis, 2023). Specifically, it has been shown that a modest rate of incorrect labels can degrade parity function learning to random guessing for bitstrings of moderate length.
@@
Building on these theoretical considerations, particularly the statistical neutrality of parity functions (Thornton, 1996), we hypothesise that AlphaZero-style algorithms, without specialised modifications, will struggle to master Nim and other impartial games on sufficiently large boards. This challenge primarily stems from the difficulty neural networks face in modelling parity functions, rather than solely from algorithmic limitations. Overcoming this requires either explicit incorporation of combinatorial game theory principles or the development of more robust neural network architectures capable of discerning such complex parity-based game properties.
@@
To explore this hypothesis and gain deeper insights into the challenges faced by RL agents in mastering Nim-like games, we developed and extensively analysed an AlphaZero-style algorithm tailored for Nim. Our investigation also involved revisiting AlphaZero’s performance in chess, focusing on positions where parity-related problems are crucial for identifying optimal moves. This comprehensive exploration culminated in a detailed theoretical and practical analysis of issues related to impartial games and parity problems. We introduce two novel metrics for evaluating self-play RL agents’ performance, ’champion’ and ’expert’ perspectives, which offer a nuanced assessment of mastery beyond conventional winning rates.
@@
Our findings reveal that AlphaZero-style algorithms exhibit significant difficulties in effectively learning to play Nim, especially as the board size increases. Despite Nim’s low inherent complexity and simple optimal strategy from both human and computational perspectives, the algorithms struggle. While the algorithm can achieve proficiency on smaller boards or towards endgames through exhaustive search, larger, non-trivial board positions frequently resulted in the policy network failing to provide valuable guidance to the Monte Carlo Tree Search (MCTS), and the value network performing no better than random guessing. These scenarios represent ’blind spots’ where optimal actions are entirely overlooked (Lan et al., 2022; Wu, 2019).
@@
In games like chess, AlphaZero’s Policy-Value Neural Networks (PV-NNs) are imperfect, but this is often compensated for by the MCTS. For some Nim boards, this imperfection might not be critical, as the policy network can guide the winning player towards positions that do not demand precise evaluations. However, we contend that this advantage diminishes on sufficiently large Nim boards due to the PV-NNs’ inherent inability to model parity functions, leading to two critical issues:
@@
The remainder of this paper is structured as follows: Section 2 introduces the Nim game and its relationship to other impartial games via the Sprague-Grundy theorem. We also discuss the theoretical limitations of neural network architectures in learning the parity function. In Section 3, we revisit AlphaZero and Leela Chess Zero (LCZero) to analyse their strengths and subtle weaknesses, particularly regarding parity-related issues impacting their PV-NN. Section 4 defines our proposed ’champion’ and ’expert’ mastery levels with illustrative examples. Section 5 presents an overview of the AlphaZero-style algorithms and our re-implementation [^1] The code for the experiments in this paper is publicly available at: https://github.com/sagebei/Impartial-games-RL. We analyse the performance of the trained algorithm from the champion and expert perspectives, respectively, and then examine the moves made on some Nim board positions. We further conducted controlled experiments to isolate whether the challenges of AlphaZero-style reinforcement learning on impartial Nim are fundamentally caused by limitations of reinforcement learning or the inherent difficulty neural networks face in learning the parity-based winning strategy required for optimal play. We conclude the paper with general remarks, conjectures and directions for further research in Section 6.
2 Impartial games and learning parity
2.1 Impartial games and Nim
@@
The class of impartial games constitutes a significant subclass of combinatorial games (Berlekamp et al., 2001, 2002, 2003, 2004). This category encompasses a diverse range of games, including classic examples such as Take-and-break, Subtraction, Heap, and Poset games. It also includes mathematically intriguing games like Nim, Sprouts, Treblecross, Cutcake, Guiles, Wyt Queens, Kayles, Grundy’s Game, Quarto, Cram, Chomp, Subtract a Square, and Notakto (Berlekamp et al., 2001, 2002). Many of these impartial games feature multiple variants, adding to their complexity and research interest.
@@
An impartial game is a two-player game characterised by identical sets of available moves from any given position, irrespective of whose turn it is. Players take alternating turns. A player loses if they cannot make a legal move on their turn, meaning the player who makes the last legal move wins. Critically, any position within an impartial game can be definitively classified as either a losing (P-position, previous player winning) or winning (N-position, next player winning) position, indicating whether the current player has no winning move or at least one winning move, respectively.
@@
Nim serves as an important example of an impartial game and is often studied within mathematical game theory due to its elegant and analytically derived solution. Nim is played by two players who take turns removing counters from distinct heaps (or piles) arranged in a row (Bouton, 1901; Nowakowski, 1998). On each turn, a player must remove at least one counter and may remove any number of counters, provided they are all drawn from a single heap. The objective is to be the player who removes the last counter, thereby leaving an empty board for the opponent. An initial Nim board can be represented as an array of heap sizes: $n_{1},n_{2},\ldots,n_{k}]$, where each $n_{k}\in\{1,3,\dots,2k-1\}$ typically denotes the initial size of the $k$-th heap. The maximum number of legal moves from an initial position $n_{1},n_{2},\ldots,n_{k}]$ is $\sum_{j=1}^{k}n_{j}$, as each move requires removing at least one counter. A game position can be represented as an array of numbers $v_{1},v_{2},\ldots,v_{k}]$, where $v_{j}\leq n_{j}$. Although Nim is often studied in abstract terms with arbitrary configurations, RL algorithms require a fixed board size, and each self-play game begins from a predefined initial state. Figure 1 visually illustrates an initial Nim board, an intermediate game state, and the terminal winning condition.
@@
Figure 1: An illustration of Nim game states. The left panel shows an initial board configuration of five heaps: $n_{1},n_{2},n_{3},n_{4},n_{5}]=[1,3,5,7,9]$. The middle panel depicts an intermediate board state during gameplay: $v_{1},v_{2},v_{3},v_{4},v_{5}]=[1,2,4,4,3]$, resulting from players removing counters. The right panel represents the terminal state where all counters are cleared, signifying a win for the player who made the last move.
@@
A property of Nim is that for any given position, it is mathematically straightforward to determine the winning player and identify all available winning moves. This determination is achieved by calculating the binary digital sum of the number of counters in each heap, a process known as the nim-sum (Bouton, 1901). This operation involves summing the heap sizes in binary without carrying over bits. The computational complexity of calculating the nim-sum has been shown to be linear in the number of heaps and logarithmic in their sizes (Fraenkel, 2004), requiring only logarithmic space memory (Calabro, 2006). Consequently, determining whether a given Nim position is a winning or losing state is computationally efficient and deterministic. This analytical solvability stands in stark contrast to the challenges faced by reinforcement learning algorithms in games lacking such perfect information or straightforward winning criteria.
@@
where $N(s)$ represents the set of all states $s^{\prime}$ reachable from $s$ in one legal move, and the mex (minimum excluded value) function returns the smallest non-negative integer not present in a given set (Beling and Rogalski, 2020). A position with Sprague-Grundy value $G(s)$ behaves identically to a Nim heap of size $G(s)$. Crucially, a position is a losing position if and only if its Sprague-Grundy value is 0. This theorem implies that the insights gained from studying Nim’s learnability extend naturally to a wide array of other impartial games.
The theoretical underpinnings of impartial games are closely connected to the nim-sum, which in turn relies on the concept of the parity function. Thus, the parity function plays an implicit or explicit central role in the theory of impartial games, as many such games can be shown to mimic Nim or parts of Nim. To illustrate this fundamental connection, consider the impartial game Sprouts (Berlekamp et al., 2001; Gardner, 1967), a relatively simple game invented by John Conway and Michael Paterson. Positions in Sprouts typically yield nimber values of 0, 1, 2, or 3 (Berlekamp et al., 2003). As depicted in Figure 2(a), a specific Sprouts position can have a nimber value of 3. Furthermore, a modified Sprouts position (Figure 2(b)), designed as an "isolated land" or "gadget" that cannot interact with external elements, also possesses a nimber value of 3. A Sprouts starting position comprising $n$ copies of the gadget shown in Figure 2(b) effectively mimics any Nim position reachable from a starting position with $n$ heaps, each containing 3 counters.
@@
(b) A Sprouts position designed as an isolated gadget, also having a nimber value of 3. This gadget can be used to construct larger impartial games that behave like Nim (adapted from p. 599 in Berlekamp et al. (2003)).
@@
Figure 2: Mimicking Nim with Sprouts gadgets. A game of Nim played on a board $3,3,\ldots,3]$ with $n$ heaps is equivalent to a Sprouts game constructed with $n$ copies of the isolated gadget shown in Figure 2(b).
@@
Despite the analytical simplicity of Nim, some impartial games present significant computational challenges, as their complexity can rise to PSPACE-complete. From an algorithmic perspective, while many impartial games share the linear-time solvability of Nim, certain others, such as Node Kayles and Geography, are demonstrably harder, being PSPACE-complete (Schaefer, 1978). Moreover, even some partisan games, where move legality depends on the player, can exhibit properties akin to impartial games, such as Dawson’s chess, which is effectively an impartial game in disguise (Berlekamp et al., 2001).
@@
Historically, algorithms for impartial games have predominantly relied on handcrafted programs incorporating principles similar to alpha-beta search, but specifically tailored for the mex operation and Sprague-Grundy values (VIENNOT, 2007). For instance, Beling and Rogalski (2020) introduced a novel method for pruning search trees based on node values computed via the mex function, particularly effective for relatively "short" impartial games like Nim, Chomp, and Cram. However, this approach typically does not scale efficiently to games with large board sizes or extensive state spaces. These conventional, handcrafted programs, which leverage the explicit mathematical structure of impartial games, serve as crucial benchmarks against which the performance and generalisation capabilities of emerging reinforcement learning-based algorithms can be rigorously evaluated. The objective is to assess whether contemporary RL methods can not only match but potentially surpass the limitations of these specialised traditional approaches, particularly for larger or more complex instances of impartial games where an explicit solution might be intractable.
2.2 Learning parity with neural networks
The parity function is a foundational challenge that reveals the inherent limitations of various neural network architectures in learning abstract, non-associative logic, a critical skill for optimal play in impartial games. Given a bitstring $x=(x_{1},x_{2},\dots,x_{n})\in\{0,1\}^{n}$, the parity function $\mathcal{P}_{n}(x)$ is formally defined as:
$\mathcal{P}_{n}(x)=\left(\sum_{i=1}^{n}x_{i}\right)\pmod{2}$ (2)
The output is $1$ (odd parity) if the number of $1$’s in the bitstring is odd, and 0 (even parity) otherwise.
Early work focused on constructing specialised NNs with fixed weights to solve the parity problem perfectly (Hohil et al., 1999; Liu et al., 2002; Franco and Cannas, 2001; Wilamowski et al., 2003). However, these models cannot generalise to real-world scenarios where the underlying data distribution is unknown and weights must be learned through training.
Focusing on the challenge of learning, Recurrent Neural Networks (RNNs), including LSTMs, face significant theoretical and practical difficulties. Although RNNs are theoretically capable of simulating any Turing Machine (Siegelmann and Sontag, 1995), this relies on unrealistic assumptions like infinite precision and computation time. In practice, training these networks is severely hampered by the vanishing gradient problem (Hochreiter and Schmidhuber, 1997).
Furthermore, the difficulty of modelling the parity function drastically escalates with bitstring length (Hahn, 2020). Since flipping a single bit alters the parity, processing long bitstrings requires persistent and perfect memorisation. RNNs and LSTMs are argued to be fundamentally ill-suited for this, as tiny memory errors can lead to disastrous consequences (Zhao et al., 2020). While parity is naturally permutation invariant, and RNNs can be regularised towards this property to improve performance (Cohen-Karlik et al., 2020), a fundamental representational bottleneck remains.
Modern architectures are similarly challenged. Self-Attention networks (Transformers (Vaswani et al., 2017)) have been shown to have strong theoretical limitations in evaluating logical formulas (like parity). Asymptotic analysis suggests that they will fail to model the parity task correctly when the input sequence is sufficiently long (Hahn, 2020).
While the adaptive computation approach of PonderNet (Banino et al., 2021) achieved near-perfect accuracy, this success highlights the increasing resource demand. It suggests that as inputs become more complex, the required computational resources become enormous (Abbe et al., 2023), and crucially, the model lacks the ability to extrapolate to input lengths beyond those seen during training.